# The role of magnetic fields in shaping $\gamma$-ray emission from the Fermi bubbles

### v1

@@ 3.1 CR source
In order to match observations which indicate that the central outflow is collimated, CRs are injected at the inner radial boundary within a half opening angle, $\theta_{1/2}=20^{\circ}$, measured from the $z$-axis. This value was motivated by observations of the Fermi bubbles and their central chimneys, i.e., a structure that appears to funnel hot gas, dust and CRs into the Galactic halo, present at the Galactic centre. For the Fermi bubbles, $\theta_{1/2}\approx 18^{\circ}-30^{\circ}$, given that the size of the Fermi bubbles is estimated to be 810$\sim$ 8~-~10∼ 8 - 10 kpc in height and 34$\sim$ 3-4∼ 3 - 4 kpc for the half-width (Carretti et al., 2013; Ackermann et al., 2014). An X-ray chimney at the base of the Fermi bubbles has been observed with a half-width of 50$\sim$ 50∼ 50 pc and a height of 150$\sim$ 150∼ 150 pc (Ponti et al., 2019). Similarly, a chimney was observed in radio emission with a half-width of 70$\sim$ 70∼ 70 pc and a height of 215$\sim$ 215∼ 215 pc (Heywood et al., 2019). Assuming a conic angle for the distribution, those two structures have $\theta_{1/2}\approx 20^{\circ}$.
@@
$Q=\frac{d\dot{N}}{dp}=\frac{\dot{N}_{\rm min}}{p_{\rm min}}\left(\frac{p}{p_{\rm min}}\right)^{-\eta},$
@@
3.2 Diffusion and advection
The distance that a CR can travel, on average, until its momentum vector is significantly changed (by $\sim$ 1 rad. in angle) by scattering off a magnetic field is determined by the diffusion tensor. In Tourmente et al. (2023), an isotropic and homogenous diffusion coefficient was adopted. Given that the MHD simulation provides the distribution of the magnetic field, here we explore the effect of adopting an anisotropic and inhomogeneous diffusion coefficient.
@@ 3.2 Diffusion and advection
The expression for the parallel diffusion scattering length, $3D_{\parallel}/c$ is given by (Jokipii, 1966; Schlickeiser, 1989),
@@
$\frac{D_{\parallel}}{c}=\frac{B^{2}_{c}}{\delta B^{2}}\left(\frac{r_{L}}{\lambda_{\mathrm{max}}}\right)^{2-\zeta}\lambda_{\mathrm{max}},$ (8)
@@
where $\zeta=5/3$ represents a Kolmogorov-type turbulence spectrum. $\lambda_{\mathrm{max}}$ represents the longest wavelength mode in the turbulence, for which we adopt $\lambda_{\rm max}=$1.5 pc. $r_{L}$ is the CR Larmor radius. $B_{c}^{2}$ is proportional to the energy density in the coherent magnetic field, and $\delta B^{2}$ is proportional to the energy density in the turbulent field. Throughout this work the ratio of the energy density in these two components is set to unity, i.e., $B_{c}^{2}=\delta B^{2}$.
The magnetic field distribution from the MHD simulations (section 2.2) has been used to define the diffusion tensor. Fig. 3 shows the 2D distribution of the parallel diffusion scattering length defined by Eq. (8) as a function of $R$ and $z$. The contour lines (dashed black line) have been drawn for a diffusion scattering length of $3D_{\parallel}(30~{\rm GeV})/c=0.3$ pc, $3D_{\parallel}(30~{\rm GeV})/c=0.42$ pc and $3D_{\parallel}(30~{\rm GeV})/c=0.9$ pc. Since the parallel diffusion scattering length depends on the magnetic field distribution, its profile can be easily understood with Figs. 2 and 2. As the radial magnetic field component $B_{r}$ dominates over $B_{\phi}$ centrally in the simulation and decreases in strength as $\propto r^{-2}$, the parallel diffusion scattering length increases as the radius increases. Beyond a radius of $\sim$ 14 kpc, $B_{\phi}$ forms a thin shell which alters how the parallel diffusion scattering length changes with distance. Beyond a radius of $r>$17 kpc, a minimal (floor) value for the magnetic field has been set to $2.6\times 10^{-5}$ $\mu$G, corresponding to $3D_{\parallel}(30~\mathrm{GeV})/c=3$ pc.
@@
Within each cell the dominant transport process is governed by the competition between the advection and diffusion rates in that cell. Fig. 3 shows the spatial distribution of the ratio between the diffusion time, $t_{\mathrm{D}}=(\Delta r)^{2}/(D_{\parallel})=(\Delta\ln r)^{2}r^{2}/(D_{\parallel})$, and the advection time, $t_{\mathrm{v}}=\Delta r/v=(\Delta\ln r)r/v$, where $\Delta\ln r$ is constant due to the logarithmic spatial grid adopted. A contour line (solid black line) has been drawn on this figure indicating when the ratio of these timescales is unity, i.e., where the diffusion time equals the advection time. Due to the small value of the total magnetic field strength (see Fig. 2), the advection time is shorter than the diffusion time in the region formed by the outflow (between 0.3 kpc and $\sim$16 kpc). This suggests that advection is the dominant CR transport mechanism. However, the magnetic field distribution still imprints itself on the results through its impact on the anisotropic CR diffusion.
@@ 3.3 Energy loss time scale and $\gamma$-ray emission
where $\kappa\approx 0.5$ represents the inelasticity of the CR in the $pp$ collision, $\sigma_{\mathrm{pp}}\approx 4\times 10^{-26}$ cm2 is the total inelastic cross-section for $pp$ collisions, and $n_{\rm gas}$ is the target gas number density (section 2.2).
[…]
For an observer, the sum of all the $\gamma$-ray emission produced along a solid angle of angular size $\Delta\Omega$, results in the total $\gamma$-ray brightness, $E_{\gamma}F_{\gamma}$. This $\gamma$-ray brightness is expressed by
[…]
@@ 4.1 CR density map
The CR transport simulations were run until the CR distribution reached a steady-state up to a Galactic height of 10$\sim$ 10∼ 10 kpc. The transport of CRs has been simulated based on the velocity and magnetic field distribution obtained from the Galactic breeze model (see section 2). Fig. 4 shows a cross-sectional slice of the resultant late-time CR density distribution, plotted in cylindrical coordinates. The CR density distribution ranges from $10^{-12}$ cm-3 (dark blue) to $10^{-9}$ cm-3 (dark red). Three contour lines are shown in Fig. 4 (dashed black lines), representing three different CR density values, $n_{\mathrm{CR}}=0.3,1$, and $3\times 10^{-11}$ cm-3, respectively.
Figure 5: The 1-3 GeV $\gamma$-ray emission map produced through the energy losses of 10-30 GeV CR. The map uses a Mollweide projection. The colour bar is linear, and the Galactic plane is masked with $E_{\gamma}F_{\gamma}=0$ at $|b|<10^{\circ}$. Both colours and mask provide a setup similar to the observational paper (Ackermann et al., 2012, Fig. 30). The resulting emission produces a bubble shape for one half of the hemisphere that appears consistent with observations, i.e., a height of $\sim 50^{\circ}$ and a width of $\sim 40^{\circ}$.
@@
The CR density distribution obtained for the case of anisotropic diffusion can be compared with that obtained for the case of isotropic diffusion shown in appendix E (see Fig. 8). As the CR diffusion is not influenced by the magnetic field direction for the isotropic diffusion case, the density distribution is more spherical in shape than for the case of anisotropic diffusion. Consequently, the CR contour lines show that the density distribution reaches a shorter distance along the $z$-axis and extends further along the $R$-axis for the isotropic diffusion case. For both anisotropic and isotropic diffusion cases considered here, CRs are injected with $\theta_{1/2}=20^{\circ}$.
@@ 4.2 $\gamma$-ray emission map
Using the obtained CR density spatial distribution, the subsequent $\gamma$-ray energy flux distribution can be determined (see section 3.3). Fig. 5 presents the skymap of the $\gamma$-ray energy flux of 1-3 GeV photons produced through inelastic $pp$ collisions of 10-30 GeV CRs with the ambient Galactic halo gas. The skymap uses Galactic coordinates where the longitude, $l$, ranges from $0^{\circ}$ to $360^{\circ}$ and the latitude, $b$, ranges from $-90^{\circ}$ to $90^{\circ}$. The colour bar range has been chosen to be similar to the $\gamma$-ray emission maps presented in Ackermann et al. (2014). For $|b|<10^{\circ}$ a mask has been applied for which $E_{\gamma}F_{\gamma}=0$, similar to Fig. 30 from Ackermann et al. (2014). For $L_{\mathrm{CR}}=1.3\times 10^{40}$ erg s-1 (with $D_{\perp}=0.1\times D_{\parallel}$ and $\theta_{\mathrm{1/2}}=20^{\circ}$), the Galactic breeze model is found to produce bilobal $\gamma$-ray emission whose brightness is comparable to that observed from the Fermi bubbles.
@@
The analysis of the energy flux distribution is taken further in Fig. 6. This figure shows the $\gamma$-ray energy flux for $40^{\circ}<|b|<50^{\circ}$ as a function of longitude for $-60^{\circ}<l<60^{\circ}$. The solid black line represents the energy flux distribution obtained from our CR transport simulation. The dotted red and blue lines, error bars and red and blue shaded regions, for which the red corresponds to the North Fermi bubbles and the blue to the South Fermi bubbles, are obtained from (Ackermann et al., 2014, Fig. 23). The lines and error bars shown in the figure correspond to the energy flux calculated using templates generated with the CR propagation code GALPROP. The shaded regions are computed for different configurations for the GALPROP template and local template proposed by the Fermi-LAT collaboration. In appendix D, Fig. 7 presents the energy flux distribution for $30^{\circ}<|b|<40^{\circ}$.
The CR injection luminosity in the simulations was set to a value of $L_{\mathrm{CR}}=1.3\times 10^{40}$ erg s-1 in order for the corresponding value of the $\gamma$-ray brightness $E_{\gamma}F_{\gamma}$ to approximately match that of the observed Fermi bubbles (see the dotted red line in Fig. 6) at $l=0^{\circ}$, $40^{\circ}<b<50^{\circ}$. For $E_{\gamma}=1~-~3$ GeV, the total corresponding $\gamma$-ray luminosity produced by the bubbles is $L_{\gamma}=2.7\times 10^{37}$ erg s-1. The simulated $\gamma$-ray brightness is consistent with the shaded red/blue regions, decreasing sharply between $0^{\circ}$ and $25^{\circ}$ of longitude, unlike the hydrodynamic model, which did not reproduce such a rapid decrease (see Fig. 5 from Tourmente et al., 2023). Given the symmetry of our simulation setup, no north-south or east-west asymmetry is produced by our model.
Once again, it is interesting to compare the results of the anisotropic diffusion case considered with those produced for the case of isotropic diffusion. In appendix E, Fig. 10 shows the $\gamma$-ray energy flux distribution for the case of isotropic diffusion. It is observed that in the case of isotropic diffusion, a more spherical shape for the $\gamma$-ray energy flux distribution is produced (see Fig. 9), with the sharp edges of the Fermi bubbles not reproduced for the case of isotropic diffusion.
Furthermore, by comparing our results for the case of the collimated injection and isotropic diffusion (see appendix E) with the results for non-collimated injection and isotropic diffusion obtained in Tourmente et al. (2023), the effect of CR collimation on the resulting $\gamma$-ray energy flux distribution can be appreciated. From Fig. 10 and Fig. 5 from Tourmente et al. (2023) one observes that the $\gamma$-ray energy flux for the case of collimated injection is more elongated in the $l$ direction and narrower in the $b$ direction, as expected.

### v2

@@ 3.1 CR source
In order to match with observations that suggest the central outflow responsible for the formation of the Fermi bubbles is collimated, we inject CRs in our simulations at the inner radial boundary within a half opening angle, $\theta_{1/2}=20^{\circ}$, measured from the $z$-axis. This value was motivated by observations of the Fermi bubbles and their central chimneys, i.e., a structure that appears to funnel hot gas, dust and CRs into the Galactic halo, present at the Galactic centre. For the Fermi bubbles, $\theta_{1/2}\approx 18^{\circ}-30^{\circ}$, given that the size of the Fermi bubbles is estimated to be $\sim 8~-~10$ kpc in height and $\sim 3-4$ kpc for the half-width (Carretti et al., 2013; Ackermann et al., 2014). An X-ray chimney at the base of the Fermi bubbles has been observed with a half-width of $\sim 50$ pc and a height of $\sim 150$ pc (Ponti et al., 2019). Similarly, a chimney was observed in radio emission with a half-width of $\sim 70$ pc and a height of $\sim 215$ pc (Heywood et al., 2019). Assuming a conic angle for the distribution, those two structures have $\theta_{1/2}\approx 20^{\circ}$. The interaction of the outflowing gas with the disc density material, although not included in our simulations (see Section 2.1.1), may be the agent responsible for the collimation of the outflow. However, we leave an investigation into this to future work.
@@
$Q=\frac{d\dot{N}}{dp}=\frac{\dot{N}_{\rm min}}{p_{\rm min}}\left(\frac{p}{p_{\rm min}}\right)^{-\eta},$ (7)
where $p_{\rm min}=1.876$ GeV/$c$, $\eta=2$ (representative of diffusive shock acceleration Bell, 1978). $\dot{N}_{\rm min}$ is a constant of normalisation representing the number of particles injected per second with momentum $p_{\rm min}$. The value of $\dot{N}_{\rm min}$ sets the CR luminosity value. For this work, $L_{\mathrm{CR}}=1.3\times 10^{40}$ erg s-1 has been chosen after comparing the produced $\gamma$-ray emission with observations (see Section 4.2). Given that it is assumed here that CRs originate from the Galactic centre, CRs are injected into the region $10~\mathrm{pc}<r<315~\mathrm{pc}$.
3.2 Diffusion and advection
@@
$\mbox{$D$}=D_{\perp}\left(\mbox{$I$}-\mbox{$\hat{b}$}\mbox{$\hat{b}$}\right)+D_{\parallel}\mbox{$\hat{b}$}\mbox{$\hat{b}$}$. (8)
The parallel diffusion coefficient, $D_{\parallel}$, represents the diffusion along the magnetic field unit vector direction, $\mbox{$\hat{b}$}\equiv\mbox{$B$}/|\mbox{$B$}|$. The perpendicular diffusion coefficient, $D_{\perp}$, represents the diffusion in the plane perpendicular to the magnetic field direction. Eq. 8 is discussed in further detail in Appendix D. Given that the MHD simulation provides the distribution of the magnetic field, we explore the effect of adopting an anisotropic and inhomogeneous diffusion coefficient.
@@
From quasi-linear theory (DeMarco et al., 2007), the parallel diffusion scattering length, $3D_{\parallel}/c$ is given by (Jokipii, 1966; Schlickeiser, 1989),
@@
$\frac{D_{\parallel}}{c}=\frac{B^{2}_{c}}{\delta B^{2}}\left(\frac{r_{L}}{\lambda_{\mathrm{max}}}\right)^{1-\zeta}r_{L},$ (9)
@@
where $\zeta=5/3$ represents a Kolmogorov-type turbulence spectrum. $\lambda_{\mathrm{max}}$ represents the longest wavelength mode in the turbulence, for which we adopt $\lambda_{\rm max}=$1.5 pc. $r_{L}$ is the CR Larmor radius. $\delta B^{2}/B_{c}^{2}$ describes the ratio of the energy density in the turbulent magnetic field to that in the coherent magnetic field. The ratio of the energy density is set to unity, i.e., $B_{c}^{2}=\delta B^{2}$.
Similarly, from quasi-linear theory (DeMarco et al., 2007), the perpendicular diffusion coefficient is given by,
$\frac{D_{\perp}}{c}=\frac{\delta B^{2}}{B_{c}^{2}}\left(\frac{r_{L}}{\lambda_{\rm max}}\right)^{\zeta-1}r_{L}.$ (10)
However, the validity of quasi-linear theory for describing the perpendicular diffusion coefficient is particularly questionable (Giacalone and Jokipii, 1999; Casse et al., 2001; Cohet and Marcowith, 2016), for which field line wandering is suggested to also play a key role in particle transport (Mertsch, 2020). We therefore instead appeal to results from approaches which attempt to go beyond this description as a motivation for the value of $D_{\perp}/D_{\parallel}$ we adopt.
Several works have provided a ratio between the perpendicular and parallel diffusion scattering lengths $\sim 10^{-2}-10^{-1}$, either directly through numerical simulations (Casse et al., 2001; Hussein et al., 2015; Cohet and Marcowith, 2016), or by attempts to derive new analytical forms for the scattering rates (Shalchi et al., 2010; Gammon and Shalchi, 2017). Here, as a first step, we arbitrarily select a ratio of $D_{\perp}=0.1D_{\parallel}$, which sits within this motivated range.
The magnetic field from the MHD simulations (section 2.2) was used to define the diffusion tensor orientation in each simulation cell. Since the turbulent magnetic field is random, and coherent on considerably smaller scales than the simulation cells, its contribution to $\hat{b}$ was neglected.
[…]
@@
Within each cell the dominant transport process is governed by the competition between the advection and diffusion rates in that cell. Fig. 3 shows the spatial distribution of the ratio between the diffusion time, $t_{\mathrm{D}}=(\Delta r)^{2}/(D_{\parallel})=(\Delta\ln r)^{2}r^{2}/(D_{\parallel})$, and the advection time, $t_{\mathrm{v}}=\Delta r/v=(\Delta\ln r)r/v$, where $\Delta\ln r$ is constant due to the logarithmic spatial grid adopted. A contour line (solid black line) has been drawn on this figure indicating when the ratio of these timescales is unity, i.e., where the diffusion time equals the advection time. Due to the small value of the total magnetic field strength (see Fig. 2), the advection time is shorter than the diffusion time in the region formed by the outflow (between 0.3 kpc and $\sim$16 kpc). This suggests that advection is the dominant CR transport mechanism. However, the magnetic field distribution still imprints itself on the results through its impact on anisotropic CR diffusion.
@@ 3.3 Energy loss time scale and $\gamma$-ray emission
where $\kappa\approx 0.5$ represents the inelasticity of the CR in the $pp$ collision, $\sigma_{\mathrm{pp}}\approx 4\times 10^{-26}$ cm2 is the total inelastic cross-section for $pp$ collisions, and $n_{\rm gas}$ is the target gas number density (section 2.2). It should be noted that the value for $\tau_{\mathrm{loss}}$ is larger than the simulation time for most CR on the simulation grid due the low density of the hot Galactic halo gas. The value for the loss time, however, imprints itself on the brightness of subsequent $\gamma$-ray emission produced (see Eq. 12).
Through inelastic CR collisions with ambient gas, approximately $\sim 10~\%$ of the relativistic proton’s energy goes into each produced $\gamma$-ray photon (Kelner et al., 2006). As the CRs considered here have an energy range of 10 to 30 GeV, photons with an energy range of $\sim 1-3$ GeV are therefore produced. The sum of all CR energy losses within the background gas gives the total non-thermal $\gamma$-ray luminosity, $L_{\gamma}$ produced, which in turn is set by the CR injection rate onto the grid, represented by $L_{\mathrm{CR}}$.
For an observer at Earth, the sum of all the $\gamma$-ray emission produced along a solid angle of angular size $\Delta\Omega$, results in the total $\gamma$-ray brightness, $E_{\gamma}F_{\gamma}$. This $\gamma$-ray brightness is expressed by
$E_{\gamma}F_{\gamma}=\frac{1}{3\Delta\Omega\tau_{\mathrm{loss}}}\int\frac{e_{\mathrm{CR}}}{r^{2}_{\mathrm{obs}}}dV=\frac{1}{12\pi\tau_{\mathrm{loss}}}\int e_{\mathrm{CR}}dr_{\mathrm{obs}},$ (12)
@@ 4.1 CR density map
The CR transport simulations were run until the CR distribution reached a steady-state up to a Galactic height of $\sim 10$ kpc, corresponding to $\sim 300$ Myr. The transport of CRs has been simulated based on the velocity and magnetic field distribution obtained from the Galactic breeze model (see Section 2). Fig. 4 shows a cross-sectional slice of the resultant late-time CR density distribution, plotted in cylindrical coordinates. The CR density distribution ranges from $10^{-12}$ cm-3 (dark blue) to $10^{-9}$ cm-3 (dark red). Three contour lines are shown in Fig. 4 (dashed black lines), representing three different CR density values, $n_{\mathrm{CR}}=0.3,1$, and $3\times 10^{-11}$ cm-3, respectively.
Figure 5: The 1-3 GeV $\gamma$-ray emission map produced through the energy losses of 10-30 GeV CR. The map uses a Mollweide projection. The colour bar is linear, and the Galactic plane is masked with $E_{\gamma}F_{\gamma}=0$ at $|b|<10^{\circ}$. Both colours and mask provide a setup similar to the observational paper (Ackermann et al., 2012, Fig. 30). The resulting emission produces a bubble shape, broadly consistent with observations (i.e. a height of $\sim 50^{\circ}$ and a width of $\sim 40^{\circ}$).
@@
The CR density distribution obtained for the case of anisotropic diffusion can be compared with that obtained for the case of isotropic diffusion shown in appendix F (see Fig. 9). As the CR diffusion is not influenced by the magnetic field direction for the isotropic diffusion case, the density distribution is more spherical in shape than for the case of anisotropic diffusion. Consequently, the CR contour lines show that the density distribution reaches a shorter distance along the $z$-axis and extends further along the $R$-axis for the isotropic diffusion case. For both anisotropic and isotropic diffusion cases considered here, CRs are injected with $\theta_{1/2}=20^{\circ}$.
@@ 4.2 $\gamma$-ray emission map
Using the obtained CR density spatial distribution, the subsequent $\gamma$-ray energy flux distribution can be determined (see Section 3.3). Fig. 5 presents the skymap of the $\gamma$-ray energy flux of 1-3 GeV photons produced through inelastic $pp$ collisions of 10-30 GeV CRs with the ambient Galactic halo gas. The skymap uses Galactic coordinates where the longitude, $l$, ranges from $0^{\circ}$ to $360^{\circ}$ and the latitude, $b$, ranges from $90^{\circ}$ to $90^{\circ}$. The colour bar range has been chosen to be similar to the $\gamma$-ray emission maps presented in Ackermann et al. (2014). For $|b|<10^{\circ}$ a mask has been applied for which $E_{\gamma}F_{\gamma}=0$, similar to Fig. 30 from Ackermann et al. (2014). For $L_{\mathrm{CR}}=1.3\times 10^{40}$ erg s-1 (with $D_{\perp}=0.1\times D_{\parallel}$ and $\theta_{\mathrm{1/2}}=20^{\circ}$), the Galactic breeze model is found to produce bilobal $\gamma$-ray emission whose brightness is comparable to that observed from the Fermi bubbles.
@@
The analysis of the energy flux distribution is taken further in Fig. 6. This figure shows the $\gamma$-ray energy flux for $40^{\circ}<|b|<50^{\circ}$ as a function of longitude for $60^{\circ}<l<60^{\circ}$. The solid black line represents the energy flux distribution obtained from our CR transport simulation. The dotted red and blue lines, error bars and red and blue shaded regions, for which the red corresponds to the North Fermi bubbles and the blue to the South Fermi bubbles, are obtained from (Ackermann et al., 2014, Fig. 23). The lines and error bars shown in the figure correspond to the energy flux calculated using templates generated with the CR propagation code GALPROP. The shaded regions are computed for different configurations for the GALPROP template and local template proposed by the Fermi-LAT collaboration. In appendix E, Fig. 8 presents the energy flux distribution for $30^{\circ}<|b|<40^{\circ}$.
The CR injection luminosity in the simulations was set to a value of $L_{\mathrm{CR}}=1.3\times 10^{40}$ erg s-1 in order for the corresponding value of the $\gamma$-ray brightness $E_{\gamma}F_{\gamma}$ to approximately match that of the observed Fermi bubbles (see the dotted red line in Fig. 6) at $l=0^{\circ}$, $40^{\circ}<b<50^{\circ}$. For $E_{\gamma}=1~-~3$ GeV, the total corresponding $\gamma$-ray luminosity produced by the bubbles is $L_{\gamma}=2.7\times 10^{37}$ erg s-1. The simulated $\gamma$-ray brightness is broadly consistent with the shaded red/blue regions, decreasing sharply between $0^{\circ}$ and $25^{\circ}$ of longitude, contrary to the hydrodynamic model, which produced a softer decrease (see Fig. 5 from Tourmente et al., 2023). Given the symmetry of our simulation setup, no north-south or east-west asymmetry is produced by our model.
Once again, it is interesting to compare the results of the anisotropic diffusion case considered with those produced for the case of isotropic diffusion. In appendix F, Fig. 11 shows the $\gamma$-ray energy flux distribution for the case of isotropic diffusion. It is observed that in the case of isotropic diffusion, a more spherical shape for the $\gamma$-ray energy flux distribution is produced (see Fig. 10), with the sharp edges of the Fermi bubbles not reproduced for the case of isotropic diffusion.
Furthermore, by comparing our results for the case of the collimated injection and isotropic diffusion (see appendix F) with the results for non-collimated injection and isotropic diffusion obtained in Tourmente et al. (2023), the effect of CR collimation on the resulting $\gamma$-ray energy flux distribution can be appreciated. From Fig. 11 and Fig. 5 from Tourmente et al. (2023) one observes that the $\gamma$-ray energy flux for the case of collimated injection is more elongated in the $l$ direction and narrower in the $b$ direction, as expected.